OpenStudy (frostypaz1):

complete the square. z^2-4z

OpenStudy (anonymous):

z^2-4z =(z-2)^2 - 4

OpenStudy (mathmate):

Hint: As an \(example\): x^2-6x =(x^2-2(3x)+3^2) -3^2 ....... create perfect square inside of brackets =(x-3)^2-3^2 =(x-3)^2-9 The objective is to make a perfect that includes all the variables, and add a constant...if necessary.

OpenStudy (frostypaz1):

Lemaz that doesnt make much sense.. I think the answer is z^2-4z+4 but im not sure

OpenStudy (mathmate):

Usually completing the square does not change the value of the expression. In your case, perhaps you are learning what a perfect square looks like, that's why you're expected to add a four to the given expression. The same term out of context can mean different things. By the way, I hope you noticed that I gave an example, it was not a solution to the problem.

OpenStudy (frostypaz1):

yeah i noticed haha